The density of a solution of potassium dichromate is 1.0665 g/cc, and it is 9.50 percent by mass. What volume of the solution (in milliliters) do you need to supply 44.0 g of potassium dichromate? please show you work
Answer #1

w/v % = w/W % x d

w/w = percent by mass = 9.50%

w/v% =percent mass by volume =?

d= density of solution = 1.0665 g/ml

w/v% = 9.50 x 1.0665 = 10.1318 %

w/v% = (w/v) x100

w =mass of solute = 44.0

v= volume of solution =?

v = [w / (w/v)%] x 100 = (44.0 x100)/ 10.1318 = 434.276 ml